OpenStudy (anonymous):

can one help me simplify each of these expressions 77.) (-j+8) 1/2 80.) (6y-1)1/4 81.) (24-24y)1/5 83.)1/3y+6-2/3y 84.)-ab^2-ab^2 and the / sign is meant for fractions not division

OpenStudy (anonymous):

77. 1/2(-j+8) = 1/2j + 4 You're basically distributing 80. 1/4(6y-1) = 3/2y-1/4 81. 1/5(24-24y)=(24/5)(1-y)

OpenStudy (anonymous):

do you have any clue to what 83 & 84 is?

OpenStudy (anonymous):

Hey :) Just do note, fractions is in fact division of the numerator by denominator so don't bother 'bout stating which. 83)y/3+(6-2)/y= y/3+4/y Change the denominator by multiplying y/3 by y and 4/y by 3 Hence (y^2 +12)/3y 84)-ab^2-ab^2 Look at this another way Ignore the negative, take it out and the equation becomes -[(ab)^2 +(ab)^2] Like a+a= 2a Answer is -[2(ab)^2] :)